# ANNZ+: an enhanced photometric redshift estimation algorithm with applications on the PAU Survey

### v1

@@ 2.2 Activation functions
Here, we summarise and review several activation functions as potential replacements for the sigmoid function originally used in annz: tanh, SiLU, softplus, ReLU, Mish, and Leaky ReLU for the benefit of photometric redshift estimation. The shapes and ranges of the activation functions are depicted in figure 2 and to encapsulate the essence of activation functions, a summary can be found in table 1.
@@ 2.2.1 Logistic Sigmoid
As mentioned in 2.1, annz used the sigmoid function as its default setting; however, this activation function faces significant problems during backpropagation. The vanishing gradient problem occurs when the sigmoid activation function is used, causing the gradient to become saturated, resulting in negligible weight updates as the network’s layers go deeper [69]. On the other hand, the exploding gradient problem happens when the gradient becomes extremely large as it propagates backwards through time. These problems cause the weight to be updated disproportionately, leading the network to diverge and fail to learn anything useful [70].
Additionally, the output range of the sigmoid function (0, 1) is not centred at 0, which can lead to poor convergence [71]. Moreover, using exponential functions in the sigmoid can increase computational demands, which might slow down the overall performance of the network.
While the sigmoid activation function is useful for certain tasks like classification [72], its drawbacks of vanishing and exploding gradient issues become significant in regression tasks (as evidenced by our findings) along with the limited output range and computational overhead, making it less desirable for more complex neural network architectures and tasks. It is also less effective when used in deeper networks [52].
@@ 3 Data
We define the testing set as the set of data without redshift information, where the aim is to determine the photo-$z$s; the performance metrics in eqs. (2.19), (2.20) and (2.21) are evaluated using this set; the training set consists of available and reliable spec-$z$s information of galaxies; while the validation set is used as part of the training process to prevent overtraining. The samples are divided equally into a 1:1:1 ratio of training-validating-testing as suggested by [38]. Sections 3.1, 3.2 and 3.3 describe the sources of the data for this study as well as the training samples used in this paper. The distribution of galaxies is illustrated in figure 3.
@@ 3.3 PAUS-GALFORM Sample
Here, we utilise the synthetic dataset from the PAUS GALFORM mock catalogue [97], where the dataset was simulated to mimic the observational characteristics of PAUS using the GALFORM semi-analytical model of galaxy formation within the Planck Millennium N-body [98]. By simulating the galaxy population to produce realistic properties such as positions, luminosities, colours, and redshifts, the PAUS-GALFORM mock catalogue serves as a valuable tool for studying galaxy evolution, validating models, and gaining insights into the physical processes shaping galaxies in the Universe. The lightcone mock catalogue is built by estimating galaxy brightness and positions by interpolating smoothly between output redshift from the Planck Millennium N-body simulation, using the point at which a galaxy crosses the observer’s lightcone (see ref. [99, 100] for further details). The mock catalogue for PAUS covers approximately 100 ${\textrm{deg}}^{2}$ and is magnitude limited to $i=22.5$, covering a redshift range $20<z<$ 2.
In our case, we did not use any photo-z errors from [101], nor did we introduce any errors in the model galaxy photometry, as done by [97], as we do not want the errors to affect the performance of the ANN alone. The photometric data has been bifurcated into two distinct samples: the PAUS-GALFORM(6), hereafter referred to as the PG-06 sample, which uses six broad bands $ugr{i^{*}}z{y^{*}}$ and the PAUS-GALFORM(46), we label as the PG-46 sample, which utilises both broad bands and the 40 narrow bands. The six bands will be used to compare the performance of the activation functions along with LRG and Stripe-82 samples, while 46 bands were used to test the impact of the activation function on high dimensional inputs.
@@ 4.1 General results
Figure 4 shows our results when testing deep and wide networks for the LRG, Stripe-82, and PAUS-GALFORM samples. The 68th percentile error for tanh, ReLU, and Leaky ReLU consistently demonstrated stable trends for all the samples, whether deep (more layers) or wide (more neurons or nodes per layer). When we delve into a comparison between sigmoid, Mish, softplus, and SiLU in a deep network, it becomes apparent that the error increases significantly as we deepen the network architecture, i.e. as we add more ”hidden” layers. On the other hand, increasing the number of nodes per layer gives us a contrasting outcome, with the results turning out to be more stable than the deep networks. However, the processing time for wide networks is long. For instance, the $10_{5}$:1 architecture only took 150 minutes to run, while it took 547 minutes for the 5:50:50:1 architecture.
@@ 4.3 Time efficiency
In terms of effectiveness, ReLU and Leaky ReLU are the most friendly functions with respect to time, volume of data and computational efficiency. These simple functions can converge faster and run all of the training data (sample sizes up to 400 000) in a short amount of time. Meanwhile, networks with sigmoid and tanh activation functions converge more slowly during training, especially in deeper and wider networks. Mish, softplus, and SiLU took a much longer time to converge than other activation functions due to their complex functional forms, and unfortunately, also produced the most problematic results. In our case, we need a faster point estimator that can converge faster during optimisation. It is confirmed that sharp minima are beneficial in terms of optimisation speed and efficiency [102]. This point gives credit to the use of ReLU and Leaky ReLU activation functions in our case.
@@ 5.1 Comparison of ANNZ+ with ANNZ2
Here, annz2 version 2.3.1 was run using artificial neural networks individually optimised for best performance, and the mean value of the PDF was chosen to represent the photo-$z$ point estimate. We evaluated the performance of annz2 using the same set of samples as presented in section 3. annz2 comes with both sigmoid and tanh activation functions, thus we used annz (utilised sigmoid function) and annz+ (utilised tanh function) to be compared with annz2’s results. To this end, we considered three distinct network architectures, which are $N_{\textrm{in}}:10:10:1$ (default), $N_{\textrm{in}}:10:10:10:10:10:10:10:10:10:10:1$ (deep) and $N_{\textrm{in}}:50:50:1$ (wide). For LRG and Stripe-82 samples, there are $N_{\textrm{in}}=5$ inputs, while for PAUS samples, $N_{\textrm{in}}=6$ and $46$ inputs were used. The purpose of this investigation was to obtain insight into the impact of network architecture on the overall performance of annz2.
@@
The discrepancies between both algorithms are due to the regularisation method, backpropagation uncertainty generation method and the optimisation process in between annz and annz2. Ref. [47] showed that annz2 performed poorly in high-dimensional input when a large number of inputs is used, which aligns with our findings. The utilisation of $k$NNs by annz2 to generate errors might be the underlying reason, it looks at the nearby objects in the significantly high-dimensional $ugr{i^{*}}z{y^{*}}+40$ narrowband space and uses them to determine the associated uncertainties. We can see that the $k$NN method probably works the best when the number of inputs is small, as the curse of dimensionality may have diluted the separating power of $k$NNs to determine better uncertainties. Despite differences in architecture, it is worth noting that annz+ demonstrate a higher level of stability than their predecessors. These observations provide strong justification for updating annz to ensure that it remains relevant and competitive alongside annz2.
@@ 6 Conclusions and future work
Further investigation into whether annz+ is consistently better than annz2 is needed. As it stands, the point photo-$z$ estimate with annz+ is indeed better. However, it might be possible to implement code into annz+ to produce PDFs and see if it can compete with annz2. This could be a good project to explore in the future.

### v2

@@ 2.2 Activation functions
Activation functions play an important role in determining the transformation of inputs and outputs within each layer of a neural network. However, under certain conditions, some activation functions may cause problems that significantly affect the network’s performance and the stability of its training process. The most common problems include:
  1. 1.
    The vanishing gradient problem, which occurs when the gradients of the loss function (which are used to update the weights of the neural network during backpropagation) become very small (close to zero) as they are propagated backwards through the layers and diminish. As a result, the network has almost no gradient to update the weights and enters into a saturated region, hindering effective learning [66]. The sigmoid and tanh functions are commonly associated with this problem.
  2. 2.
    The exploding gradient problem, which arises due to poor weight initialisation or when the gradients of the loss function become extremely large as it propagates backwards through layers, causing the weight to be updated disproportionately, leading the network to diverge or produce ’NaN’ values and fail to learn anything useful [67]. Our results showed that the Softplus and Mish activation functions experience these issues when applied within complex network architectures.
  3. 3.
    The dying gradient problem, which refers to a situation where certain neurons stop learning altogether, resulting in “dead” neurons that contribute nothing to the model’s output. This typically occurs with activation functions like ReLU, which produces zero output for all negative inputs. Consequently, the gradients of the loss function become zero during backpropagation, causing the weights to remain unchanged and preventing any further learning [68].
By understanding these issues upfront, we can analyse the impact of the activation functions on the performance of annz. Here, we summarise and review several activation functions as potential replacements for the sigmoid function originally used in annz: tanh, SiLU, softplus, ReLU, Mish, and Leaky ReLU for the benefit of photometric redshift estimation. The shapes and ranges of the activation functions are depicted in figure 2 and to encapsulate the essence of activation functions, a summary can be found in table 1.
@@ 2.2.1 Logistic Sigmoid
As mentioned in 2.1, annz used the sigmoid function as its default setting; however, this activation function faces significant problems during backpropagation. The vanishing gradient problem occurs when the outputs of the sigmoid function are very close to 1 or 0, causing the function to saturate when the inputs are large, either positively or negatively. Furthermore, the derivative of the sigmoid function is capped at a maximum of 0.25 (when the input is 0) and decreases rapidly as the input moves away from 0.
Additionally, the output range of the sigmoid function (0, 1) is not centred at 0, which can lead to poor convergence [72]. Moreover, using exponential functions in the sigmoid can increase computational demands, which might slow down the overall performance of the network. While the sigmoid activation function is useful for certain tasks like classification [73], its drawbacks of vanishing and exploding gradient issues become significant in regression tasks (as evidenced by our findings) along with the limited output range and computational overhead, making it less desirable for more complex neural network architectures and tasks. It is also less effective when used in deeper networks [52].
@@ 3 Data
We define the testing set as the set of data without redshift information, where the aim is to determine the photo-$z$s; the performance metrics in eqs. (2.19), (2.20) and (2.21) are evaluated using this set; the training set consists of available and reliable spec-$z$s information of galaxies; while the validation set is used as part of the training process to prevent overtraining. The total number of galaxies in the samples is divided equally into a 1:1:1 ratio of training-validating-testing sets as suggested by [38]. Sections 3.1, 3.2 and 3.3 describe the sources of the data for this study as well as the training samples used in this paper. The distribution of galaxies is illustrated in figure 3.
@@ 3.3 PAUS-GALFORM Sample
Here, we utilise the synthetic dataset from the PAUS GALFORM mock catalogue [97], which consists of 14,053 galaxies. The dataset was simulated to mimic the observational characteristics of PAUS using the GALFORM semi-analytical model of galaxy formation within the Planck Millennium N-body [98]. By simulating the galaxy population to produce realistic properties such as positions, luminosities, colours, and redshifts, the PAUS-GALFORM mock catalogue serves as a valuable tool for studying galaxy evolution, validating models, and gaining insights into the physical processes shaping galaxies in the Universe. The lightcone mock catalogue is built by estimating galaxy brightness and positions by interpolating smoothly between output redshift from the Planck Millennium N-body simulation, using the point at which a galaxy crosses the observer’s lightcone (see ref. [99, 100] for further details). The mock catalogue for PAUS covers approximately 100 ${\textrm{deg}}^{2}$ and is magnitude limited to $i=22.5$, covering a redshift range $20<z<$ 2.
In our case, we did not use any photometric errors from [101], nor did we introduce any errors in the model galaxy photometry, as done by [97], as we do not want the errors to affect the performance of the ANN alone. However, we acknowledge that the photo-$z$ generated from the PAUS-GALFORM sample is constructed based on a limited number of initial parameters and synthetic magnitudes, this idealisation could allow annz+ to learn the underlying parameterisation of the semi-analytic model. This is unlike observed data such as the SDSS LRG and Stripe-82 samples, which have added real-world variability. By leveraging both mock and observed datasets, we provide a balanced assessment that highlights the robustness of the activation functions across different data scenarios and ensure that the conclusions drawn from the mock data are not overly dependent on its idealised characteristics. The photometric data has been bifurcated into two distinct samples: the PAUS-GALFORM(6), hereafter referred to as the PG-06 sample, which uses six broad bands $ugr{i^{*}}z{y^{*}}$ and the PAUS-GALFORM(46), we label as the PG-46 sample, which utilises both broad bands and the 40 narrow bands. The six bands will be used to compare the performance of the activation functions along with LRG and Stripe-82 samples, while 46 bands were used to test the impact of the activation function on high dimensional inputs.
@@ 4.1 General results
Figure 4 shows our results when testing deep and wide networks for the LRG, Stripe-82, and PAUS-GALFORM samples. The 68th percentile error for tanh, ReLU, and Leaky ReLU consistently demonstrated stable trends for all the samples, whether deep (more layers) or wide (more neurons or nodes per layer). When we delve into a comparison between sigmoid, Mish, softplus, and SiLU in a deep network, it becomes apparent that the error increases significantly as we deepen the network architecture, i.e. as we add more ”hidden” layers. On the other hand, increasing the number of nodes per layer gives us a contrasting outcome, with the results turning out to be more stable than the deep networks. However, the processing time for wide networks is long. For instance, the computational runtime required to process the largest sample used in this work (i.e. the LRG sample) was 150 minutes for the $10_{5}$:1 architecture, while it took 547 minutes for the 5:50:50:1 architecture. The LRG sample contains 412,138 objects, leading to a mean runtime per object for each architecture to be approximately 0.02 seconds and 0.08 seconds, respectively. In the context of high-dimensional input data, the PG-46 sample took a maximum runtime of 36 minutes for deep networks and 150 minutes for wide networks, utilising the mentioned architecture. The average runtime per object, across a total of 14,053 objects was 0.15 seconds (deep) and 0.64 seconds (wide). Consequently, the increased number of inputs resulted in a longer average runtime per object.
@@ 4.3 Time efficiency
Functions Runtime (minutes) $\sigma_{\textrm{RMS}}$ $\sigma_{\textrm{68}}$
sigmoid 43 0 .02811 0.02285
tanh 56 0.02812 0.02285
SiLU 32 0.02822 0.02290
Softplus 37 0.02864 0.02329
ReLU 27 0.02814 0.02283
Mish 60 0.02841 0.02296
Leaky ReLU 27 0.02814 0.02287
Table 3: Comparison of runtime for deriving photo-$z$ estimates from a LRG sample containing 412,138 galaxies, alongside the performance metrics ($\sigma_{\textrm{RMS}}$ and $\sigma_{\textrm{68}}$) for various activation functions using a 5:10:10:1 network architecture.
In terms of effectiveness, ReLU and Leaky ReLU are the most friendly functions with respect to time, volume of data and computational efficiency. According to table 3, these simple functions can converge faster and run all of the training data (sample sizes up to 400 000) in 27 minutes with competitive $\sigma_{\textrm{RMS}}$ and $\sigma_{\textrm{68}}$ values. Meanwhile, networks with sigmoid and tanh activation functions converge more slowly with respective runtimes of 43 and 56 minutes for simple architecture. A longer runtime is expected when the complexity of the network increases, especially in deeper and wider networks. Mish took the longest runtime to converge with a total time of 60 minutes due to its complex functional forms, and unfortunately, also produced the most problematic results, followed by SiLU and Softplus. In our case, we need a faster point estimator that can converge faster during optimisation. It is confirmed that sharp minima are beneficial in terms of optimisation speed and efficiency [102]. This point gives credit to the use of ReLU and Leaky ReLU activation functions for managing extensive data sets associated with large-scale surveys like Euclid and LSST, which aim to calculate photo-$z$ for billions of objects.
@@ 5.1 Comparison of ANNZ+ with ANNZ2
Here, annz2 version 2.3.1 was run using artificial neural networks individually optimised for best performance. The configurations used were NCycles=$9$, EstimatorType=MSE, TrainingMethod=BFGS, LearningRate=0.02 and UseRegulator=False. We evaluated the performance of annz2 using the same set of samples as presented in section 3. annz2 comes with both sigmoid and tanh activation functions, thus we used annz (utilised sigmoid function) and annz+ (utilised tanh function) to be compared with annz2’s results. To this end, we considered three distinct network architectures, which are $N_{\textrm{in}}:10:10:1$ (default), $N_{\textrm{in}}:10:10:10:10:10:10:10:10:10:10:1$ (deep) and $N_{\textrm{in}}:50:50:1$ (wide). For LRG and Stripe-82 samples, there are $N_{\textrm{in}}=5$ inputs, while for PAUS samples, $N_{\textrm{in}}=6$ and $46$ inputs were used.
The mean value of the PDF ($z_{\textrm{pdf}}$) was chosen to represent the photo-$z$ point estimate for annz2, which is calculated via a weighted average across all runs. This is because ref. [38] found that $z_{\textrm{pdf}}$ produces better overall accuracy at the high redshift end as compared to the peak (mode) value of the PDF ($z_{\textrm{peak}}$). However, if the PDF has a low redshift contribution, we acknowledge that $z_{\textrm{peak}}$ may represent the most suitable value that could yield different results from annz2. Moreover, regardless of the point estimate method chosen, whether the mean or mode—annz+ consistently outperforms annz2. It achieves a 3.6 per cent improvement in the LRG sample, a 9.5 per cent in the Stripe-82 sample, and a 25.9 per cent in the PG-46 sample. These comparisons remain valid for annz2 when it employs results generated from $z_{\textrm{peak}}$, utilised tanh activation function and adheres to the default architecture. The purpose of this investigation was to obtain insight into the impact of network architecture on the overall performance of annz2.
@@
The discrepancies between both algorithms are due to the regularisation method, backpropagation uncertainty generation method and the optimisation process in between annz and annz2. Ref. [47] showed that annz2 performed poorly in high-dimensional input when a large number of inputs is used, which aligns with our findings. The utilisation of $k$NNs by annz2 to generate errors might be the underlying reason, it looks at the nearby objects in the significantly high-dimensional $ugr{i^{*}}z{y^{*}}+40$ narrowband space and uses them to determine the associated uncertainties. We can see that the $k$NN method probably works the best when the number of inputs is small, as the curse of dimensionality may have diluted the separating power of $k$NNs to determine better uncertainties. Despite differences in architecture, it is worth noting that annz+ demonstrate a higher level of stability than their predecessors. We also recognise that annz2 was designed to generate accurate PDFs, thus it may not be optimised for producing derived point estimates. These observations provide strong justification for updating annz to ensure that it remains relevant and competitive alongside annz2.
@@ 6 Conclusions and future work
Further investigation into whether annz+ is consistently better than annz2 is needed. As it stands, the point photo-$z$ estimate with annz+ is indeed better. However, upcoming surveys such as Euclid and LSST will probe higher redshifts and the complexity of photometric redshift estimation will increase due to redshift degeneracies, such as the Lyman/Balmer break degeneracy, which becomes more prominent for samples spanning $32<z<$ 3. In such cases, algorithms that are capable of producing PDFs rather than point estimates are likely to perform better and thus are more valued by the community. While annz+ currently focuses on point estimates, it might be possible to modify annz+ to produce PDFs and see if it can compete with annz2. To implement PDF generation within the annz+ framework, several modifications are necessary. First, we need to write additional C++ scripts to locate the output layer configuration, allowing for the production of either Gaussian parameters or binned probabilities. Next, we should define the loss function to align with these modifications. Finally, we will need to create additional scripts to plot the PDF. This could be a good project to explore in the future.